# Personalized Federated Learning with Heat-Kernel Enhanced Tensorized Multi-View Clustering

### v1

@@ 4.4 Complete Tensorized Federated Algorithm
return Global tensor model $(\mathcal{G}_{g}^{(T)},\mathbf{P}_{g}^{(T)},\mathbf{Q}_{g}^{(T)},\mathbf{R}_{g}^{(T)},\mathbf{v}_{g}^{(T)})$ and personalized models $\{(\mathcal{G}_{i}^{(T)},\mathbf{P}_{i}^{(T)},\mathbf{Q}_{i}^{(T)},\mathbf{R}_{i}^{(T)},\mathbf{v}_{i}^{(T)})\}_{i=1}^{M}$
5 Theoretical Analysis
5.1 Convergence Properties
@@
For each client $i$, the local heat-kernel enhanced multi-view fuzzy clustering algorithm converges to a local minimum of the objective function $J_{i}^{local}$.
@@
    Membership updates (Eq. 26) minimize $J_{i}^{local}$ w.r.t. $U_{i}$ for fixed $A_{i},V_{i}$
@@
    Center updates (Eq. 27) minimize $J_{i}^{local}$ w.r.t. $A_{i}$ for fixed $U_{i},V_{i}$
@@
    View weight updates (Eq. 28) minimize $J_{i}^{local}$ w.r.t. $V_{i}$ for fixed $U_{i},A_{i}$
Since $J_{i}^{local}$ is bounded below by zero and decreases monotonically, convergence is guaranteed. ∎
@@
Under mild regularity conditions, the federated algorithm converges to a stationary point of the global objective function $J_{global}$.
@@ 5.2 Privacy Guarantees
The shared statistics $S_{i}^{centers},S_{i}^{views},S_{i}^{quality}$ have bounded sensitivity. Adding Gaussian noise with variance $\sigma^{2}=\frac{2\Delta^{2}\log(1.25/\delta)}{\epsilon^{2}}$ where $\Delta$ is the sensitivity, ensures $(\epsilon,\delta)$-differential privacy. ∎
@@
The federated protocol requires $\mathcal{O}(c\cdot\bar{d}\cdot\bar{s})$ communication per round, where $\bar{d}$ and $\bar{s}$ are average feature dimensions and views across clients.
5.3 Computational Complexity
@@
The local clustering complexity per client $i$ per iteration is $\mathcal{O}(n_{i}cs_{i}D_{i})$ where $D_{i}=\sum_{h=1}^{s_{i}}d_{h}^{(i)}$ is the total local dimensionality.
@@
5.4 Tensorized Complexity Analysis
@@
The tensorized local clustering complexity per client $i$ per iteration is $\mathcal{O}(n_{i}cr_{2}r_{3}+n_{i}cD_{i}s_{i})$ where $r_{2},r_{3}$ are the tensor decomposition ranks.
@@
    Heat-kernel coefficient computation: $\mathcal{O}(n_{i}D_{i}s_{i})$
@@
    Tensor reconstruction: $\mathcal{O}(cr_{2}r_{3}+cD_{i}r_{2}+cs_{i}r_{3})$
@@
    Membership updates: $\mathcal{O}(n_{i}cs_{i}D_{i})$
@@
    Tucker decomposition updates: $\mathcal{O}(n_{i}cr_{2}r_{3})$ per factor using ALS
The dominant term is $\mathcal{O}(n_{i}cD_{i}s_{i})$ when $r_{2},r_{3}\ll\min(D_{i},s_{i})$. ∎
@@
The tensorized federated protocol requires $\mathcal{O}(cr_{2}r_{3}+c^{2}+\bar{D}r_{2}+\bar{s}r_{3})$ communication per round, compared to $\mathcal{O}(c\bar{D}\bar{s})$ for the non-tensorized version.
The tensorized approach provides significant communication savings when $r_{2}\ll\bar{D}$ and $r_{3}\ll\bar{s}$, which is typically the case in practice.
@@
$\|\mathcal{A}_{i}-\mathcal{G}_{i}\times_{1}\mathbf{P}_{i}\times_{2}\mathbf{Q}_{i}\times_{3}\mathbf{R}_{i}\|_{F}^{2}\leq\epsilon_{tensor}$ (100)
@@
5.5 Personalization-Accuracy Trade-off
@@
$\|A_{i}-A_{g}\|_{F}^{2}\leq\frac{\lambda_{i}^{2}}{1-\lambda_{i}^{2}}\|A_{i}^{local}-A_{g}\|_{F}^{2}$ (101)
where $A_{i}^{local}$ is the purely local optimum.
This bound shows that personalization parameters $\lambda_{i}$ control the trade-off between local specialization and global coordination.
@@ 6 Conclusion
This paper presents a novel personalized federated learning framework that integrates heat-kernel enhanced multi-view clustering with tensor-based global coordination. The key contributions include:
  1. 1.
    Heat-kernel Enhanced Federated Clustering: We adapt heat-kernel coefficients to federated multi-view clustering, enabling more effective capture of complex local patterns while maintaining global coordination.
  2. 2.
    Personalized Aggregation Mechanism: Our framework provides adaptive personalization that balances local specialization with global knowledge sharing through learnable personalization parameters.
  3. 3.
    Privacy-Preserving Protocol: The aggregation protocol ensures differential privacy while maintaining clustering effectiveness through carefully designed statistical sharing.
  4. 4.
    Comprehensive Theoretical Analysis: We provide convergence guarantees, privacy bounds, and complexity analysis that establish the theoretical foundation of the approach.
@@ 6.1 Future Directions
Several promising research directions emerge from this work:
  • Dynamic View Discovery: Extending the framework to handle dynamically changing view structures across federated rounds.
  • Hierarchical Federated Clustering: Investigating multi-level federated architectures for large-scale deployment.
  • Adversarial Robustness: Analyzing and improving robustness against malicious clients and data poisoning attacks.
  • Continuous Learning: Adapting the framework for federated continual learning scenarios where data distributions evolve over time.
  • Cross-modal Applications: Exploring applications to federated learning with fundamentally different data modalities (text, images, sensors, etc.).
The proposed framework opens new possibilities for privacy-preserving, personalized machine learning on distributed multi-view data, with broad applications in healthcare, IoT, and collaborative intelligence systems.

### v2

@@ 6.6.3 Tensorized Personalized Update Mechanism
return Global tensor model $(\mathcal{G}_{\text{global}}^{(T)},\mathbf{P}_{\text{global}}^{(T)},\mathbf{Q}_{\text{global}}^{(T)},\mathbf{R}_{\text{global}}^{(T)},A_{\text{global}}^{(T)},V_{\text{global}}^{(T)})$ and personalized models $\{(\mathcal{G}_{[\ell]}^{(T)},\mathbf{P}_{[\ell]}^{(T)},\mathbf{Q}_{[\ell]}^{(T)},\mathbf{R}_{[\ell]}^{(T)},A_{[\ell]}^{(T)},V_{[\ell]}^{(T)})\}_{\ell=1}^{M}$
6.7 Theoretical Analysis
6.7.1 Convergence Properties
@@
For each client $\ell$, the local heat-kernel enhanced multi-view fuzzy clustering algorithm converges to a local minimum of the objective function $J_{[\ell]}$.
@@
    Membership updates (Eq. 26) minimize $J_{[\ell]}$ w.r.t. $U_{[\ell]}^{*}$ for fixed $A_{[\ell]},V_{[\ell]}$
@@
    Center updates (Eq. 27) minimize $J_{[\ell]}$ w.r.t. $A_{[\ell]}$ for fixed $U_{[\ell]},V_{[\ell]}$
@@
    View weight updates (Eq. 28) minimize $J_{[\ell]}$ w.r.t. $V_{[\ell]}$ for fixed $U_{[\ell]},A_{[\ell]}$
Since $J_{[\ell]}$ is bounded below by zero and decreases monotonically, convergence is guaranteed. ∎
@@
Under mild regularity conditions, the federated algorithm converges to a stationary point of the global objective function $J_{\text{global}}$.
@@ 6.7.2 Privacy Guarantees
The shared statistics $S_{[\ell]}^{centers},S_{[\ell]}^{views},S_{[\ell]}^{quality}$ have bounded sensitivity. Adding Gaussian noise with variance $\sigma^{2}=\frac{2\Delta^{2}\log(1.25/\delta)}{\epsilon^{2}}$ where $\Delta$ is the sensitivity, ensures $(\epsilon,\delta)$-differential privacy. ∎
@@
The federated protocol requires $\mathcal{O}(c(\ell)\cdot\bar{d}_{h}^{(\ell)}\cdot\bar{s}(\ell))$ communication per round, where $\bar{d}_{h}^{(\ell)}$ and $\bar{s}(\ell)$ are average feature dimensions and views across clients.
6.7.3 Computational Complexity
@@
The local clustering complexity per client $\ell$ per iteration is $\mathcal{O}(n(\ell)c(\ell)s(\ell)D(\ell))$ where $D(\ell)=\sum_{h=1}^{s(\ell)}d_{h}^{(\ell)}$ is the total local dimensionality.
@@
6.7.4 Tensorized Complexity Analysis
@@
The tensorized local clustering complexity per client $\ell$ per iteration is $\mathcal{O}(n(\ell)c(\ell)r_{2}r_{3}+n(\ell)cD(\ell)s(\ell))$ where $r_{2},r_{3}$ are the tensor decomposition ranks.
@@
    Heat-kernel coefficient computation: $\mathcal{O}(n(\ell)D(\ell)s(\ell))$
@@
    Tensor reconstruction: $\mathcal{O}(c(\ell)r_{2}r_{3}+c(\ell)D(\ell)r_{2}+c(\ell)s(\ell)r_{3})$
@@
    Membership updates: $\mathcal{O}(n(\ell)c(\ell)s(\ell)D(\ell))$
@@
    Tucker decomposition updates: $\mathcal{O}(n(\ell)c(\ell)r_{2}r_{3})$ per factor using ALS
The dominant term is $\mathcal{O}(n(\ell)c(\ell)D(\ell)s(\ell))$ when $r_{2},r_{3}\ll\min(D(\ell),s(\ell))$. ∎
@@
The tensorized federated protocol requires $\mathcal{O}(c(\ell)r_{2}r_{3}+c(\ell)^{2}+\bar{D}(\ell)r_{2}+\bar{s}(\ell)r_{3})$ communication per round, compared to $\mathcal{O}(c(\ell)\bar{D}(\ell)\bar{s}(\ell))$ for the non-tensorized version.
The tensorized approach provides significant communication savings when $r_{2}\ll\bar{D}(\ell)$ and $r_{3}\ll\bar{s}(\ell)$, which is typically the case in practice.
@@
$\|\mathcal{A}_{[\ell]}-\mathcal{G}_{[\ell]}\times_{1}\mathbf{P}_{[\ell]}\times_{2}\mathbf{Q}_{[\ell]}\times_{3}\mathbf{R}_{[\ell]}\|_{F}^{2}\leq\epsilon_{tensor}$ (111)
@@
6.7.5 Personalization-Accuracy Trade-off
@@
$\|A_{[\ell]}-A_{\text{global}}\|_{F}^{2}\leq\frac{\lambda_{\ell}^{2}}{1-\lambda_{\ell}^{2}}\|A_{[\ell]}-A_{\text{global}}\|_{F}^{2}$ (112)
where $A_{[\ell]}$ is the purely local optimum.
This bound shows that personalization parameters $\lambda_{\ell}$ control the trade-off between local specialization and global coordination.
@@ 7 Conclusion
The present paper proposes a novel personalized federated learning framework that integrates heat-kernel enhanced multi-view clustering with tensor-based global coordination. The primary contributions of this study include the introduction of a novel method, Heat-kernel Enhanced Federated Clustering, which has not been previously documented. The adoption of heat-kernel coefficients in federated multi-view clustering facilitates the more effective capture of complex local patterns while ensuring the maintenance of global coordination. In the Personalized Aggregation Mechanism, the framework provides an adaptive personalization that balances local specialization with global knowledge sharing through learnable personalization parameters. In the context of Privacy-Preserving Protocol, the aggregation protocol has been engineered to ensure differential privacy while preserving the efficacy of clustering through the implementation of meticulously designed statistical sharing methodologies. A comprehensive theoretical analysis is provided, offering convergence guarantees, privacy bounds, and complexity analysis. These elements serve to establish the theoretical foundation of the approach.
@@ 7.1 Future Directions
This work suggests several promising research directions. Initially, the Dynamic View Discovery method was employed. The framework is being extended to accommodate the dynamic alteration of view structures during federated rounds. The present study explores the potential of hierarchical federated clustering for large-scale deployment, investigating multi-level federated architectures in the process. In the context of adversarial robustness, the focus is on the analysis and enhancement of systems’ resilience to malicious clients and data poisoning attacks. The concept of continuous learning involves the adaptation of a framework for federated continual learning scenarios, with the understanding that data distributions undergo evolution over time. In the domain of cross-modal applications, the focus is on the exploration of applications to federated learning with fundamentally different data modalities, including but not limited to text, images, and sensors. The proposed framework presents novel opportunities for privacy-preserving, personalized machine learning on distributed multi-view data, with extensive applications in healthcare, the Internet of Things (IoT), and collaborative intelligence systems. The proposed framework opens new possibilities for privacy-preserving, personalized machine learning on distributed multi-view data, with broad applications in healthcare, IoT, and collaborative intelligence systems.
7.2 Potential Limitations
The proposed frameworks, while innovative, present several potential limitations. The computational complexity associated with tensor decompositions may pose challenges for clients with limited resources. The selection of personalization parameters requires careful tuning to balance local adaptation and global coordination effectively. The assumption of data homogeneity across clients may not hold in all scenarios, potentially impacting clustering performance. Additionally, the privacy guarantees depend on the proper calibration of noise, which may affect the utility of the shared statistics. Future work will focus on addressing these limitations through algorithmic optimizations, adaptive parameter selection methods, and robust privacy-preserving techniques with minimal utility loss and enhanced adaptability to heterogeneous data distributions. Validations both on synthetic and real-world datasets will be conducted to assess the practical performance and limitations of the proposed frameworks.
7.3 Applications and Implications
The proposed personalized federated heat-kernel enhanced multi-view clustering framework has significant applications across various domains. In healthcare, it can facilitate collaborative analysis of patient data from multiple institutions while preserving privacy. In IoT, the framework can enable efficient clustering of sensor data collected from distributed devices, enhancing real-time decision-making. In collaborative intelligence systems, it can support the integration of diverse data sources for improved insights and predictions. The implications of this work extend to advancing privacy-preserving machine learning techniques, promoting personalized models that respect individual client characteristics, and fostering collaboration in distributed environments without compromising data security. With the aggregation of several factors such as privacy, personalization, and multi-view data integration using tensorization, this framework sets a new standard for federated learning applications in complex, real-world scenarios.
Acknowledgment: This paper was authored during the period of the author’s postdoctoral studies at ISTI-CNR, Italy.
Funding: This research received no external funding.
Conflicts of interest: The author declare no conflict of interest.